The president of Lowell Inc. has asked you to evaluate the proposed acquisition of a new computer. The computer's price is $60,000, and it falls into the MACRS 3-year class (33% in year 1, 45% in year 2, 15% in year 3, and 7% in year 4). Purchase of the computer would require an increase in net operating working capital of $2,000. The computer would increase the firm's before-tax revenues by $20,000 per year but would also increase operating costs by $5,000 per year. The computer is expected to be used for 4 years and then be sold for $25,000. The firm's marginal tax rate is 40 percent, and the project's cost of capital is 14 percent. What is the total value of the terminal year non-operating cash flows (after-tax salvage value + working capital recovered) at the end of Year 4?

a.

$17,000

b.

$18,680

c.

$21,000

d.

$25,000

e.

$27,000
Answer #1

Please provide rating.....

Post tax salvage value for asset = (Sale price - Depreciable value at the end of 4th year)*(1-Tax rate)
Sales price = 25000
Depreciable value= 0
Tax rate = 40%
Post tax salvage value = =25000*(1-40%)
Post tax salvage value = 15000
Recovery of working capital = 2000
Therefore non operating terminal year cash flow = 15000+2000 =17000
Correct answer is option a. 17000